# Understanding Sampler Stochasticity in Training Diffusion Models for RLHF

### v1

@@ 3.1 VE with a Gaussian target
$dY_{t}^{REF}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}Y_{t}^{REF}dt+\eta\sqrt{2(T-t)}dB_{t}.$
Next we set the reward function $r(x)=-(x-1)^{2}$, so the goal of fine-tuning is to drive the sample towards the mode $1$. We also specify the fine-tuned SDE and ODE by
$dY_{t}^{SDE}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}(Y_{t}^{SDE}+\theta_{\eta}^{*})dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{ODE}=-\frac{T-t}{1+(T-t)^{2}}(Y_{t}^{ODE}+\theta_{\eta}^{*})dt,$
@@
For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-(1+T^{2})^{-\frac{1+\eta^{2}}{2}}\right)\right]^{-1}$. Moreover,
@@ 3.2 VP with a Gaussian target
Under the same setup as in the VE case, we have:
$dY_{t}^{REF}=-\eta^{2}(T-t)Y_{t}^{REF}dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{SDE}=-\eta^{2}(T-t)Y_{t}^{SDE}\,dt-(1+\eta^{2})(T-t)\theta^{*}_{\eta}(t)\,dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{ODE}=-(T-t)\theta_{\eta}^{*}(t)dt,$
@@
Consider the Variance Preserving model (equation 12), with the reward function $r(x)=-(x-1)^{2}$. For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-e^{-\frac{(1+\eta^{2})T^{2}}{2}}\right)\right]^{-1}$. Moreover,
@@ 3.3 VE/VP with a mixture Gaussian target
For all these examples, the reward gap $\Delta_{\eta}\to 0$ (independent of $\eta$), as $T\to\infty$. Such a phenomenon will also be observed in fine-tuning T2I models with more complex rewards.
@@ 4 Non-vacuous bound on the reward gap
In Section 3, we studied the reward gap for the VE and VP models with a Gaussian/mixture Gaussian target distribution. The analysis relies on the explicit computation of the score function, which is not available in general. Here our goal is to provide a bound on $W_{2}(Y_{T}^{SDE},Y_{T}^{ODE})$, where
$dY_{t}^{SDE}=\left(-f(T-t,Y_{t}^{SDE})+\frac{1+\eta^{2}}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{SDE})\right)dt+\eta\,g(T-t)dB_{t},$
$dY_{t}^{ODE}=\left(-f(T-t,Y_{t}^{ODE})+\frac{1}{2}g^{2}(T-t)s_{\theta_{\eta}^{*}}(T-t,Y_{t}^{ODE})\right)dt.$
Here we fine-tune directly on the score matching function $\{s_{\theta}(t,x)\}_{\theta}$, instead of adding a control as in Section 3. We need the following assumptions.
@@
The following conditions hold:
@@
  3. 3.
[…]
The conditions 1 (dissipativity of $f$) and 2 (Lipschitz condition on $s_{\theta}$) are standard in stability analysis. The condition 3 ensures process $\{Y^{SDE}_{t}\}$ a contractive drift, which holds if $f$ is dissipative, and the fine-tuned distribution satisfies some strongly log-concave condition, see e.g., (Gao et al., 2025; Tang & Zhao, 2024). The following theorem gives a non‑vacuous bound on the reward gap, see proof in Appendix D.
@@
[…]
@@ 5 Numerical experiments
In this section, we conduct numerical experiments for fine-tuning large-scale T2I models by leveraging sampler stochasticity, and examine the reward gap. Two classes of RL algorithms are considered: DDPO (Black et al., 2024) and MixGRPO (Li et al., 2025). The preference rewards that we use for fine-tuning DDPO include the LAION aesthetic (Schuhmann et al., 2022), HPS_v2.1 (Wu et al., 2023), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2023). The rewards for fine-tuning MixGRPO include HPS_v2.1, PickScore, ImageReward, and Unified Reward. (Wang et al., 2025) To study the reward gap, we use a high stochasticity $\eta$ under gDDIM scheme to generate training samples, and compare them with the ODE samples generated using the same weights in each iteration. The stochasticity scale $\eta$ controls the backward dynamics according to equation 4 and equation 5 in Section 2.2. Choices of hypermaraters are detailed in Appendix E.
@@ 5.1 DDPO
We use Stable Diffusion v1.5 (Rombach et al., 2022) as the base model, and fine-tune it with DDPO. During training, we adopt ImageReward and PickScore as preference rewards, while Aesthetic and HPS_v2 are included as additional evaluation metrics. Figure 1 (and Appendix G) provide representative generations from the fine-tuned models. Our main observations are summarized as follows:
$\bullet$ Decreasing Reward Gap with Quality Improvement: we conduct fine-tuning experiments under PickScore with $\eta=1.2$ and calculate SDE–ODE reward (contrary to Section 3, here we subtract ODE reward value from SDE reward for non-negativity) differences under multiple preference functions every 200 steps until reward collapse. As shown in Table 2, the gap decreases as image quality improves for both samplers.
$\bullet$ High Stochasticity Benefits Moderate Time Steps: we compare fine-tuning under ImageReward and PickScore at $\eta\in\{1.0,1.2,1.5\}$ for 200 steps. As shown in Table 1, $\eta=1.2$ under ImageReward achieves the best in-domain and out-of-domain performance, while PickScore’s performances depend on evaluation metrics.
$\bullet$ Richer Prompt Contents Reduce Reward Gap: we compare performances with animal versus more comprehensive prompts (Figure 1, Appendix G) under ImageReward with $\eta=1.2$. As shown in Table 3, complex prompts generate higher post-tuning rewards and higher in-group variance in post-training. Moreover, their richer instructions reduce the SDE–ODE reward gap.
Refer to caption Refer to caption
Figure 3: Evolution of reward gap during DDPO training under PickScore fine-tuning with stochasticity $\eta\in\{1.0,1.2,1.5\}$. The gap for multiple results decreases steadily as training progresses, indicating that image quality improves for both samplers and that ODE inference remains competitive.
$T=\bm{0}$ 200 400 600 800 1000 1200 1400 (1476)
@@
SDE Reward 20.823 21.310 21.647 21.901 22.068 22.265 22.306 22.417 (18.920)
ODE Reward 20.730 21.165 21.500 21.783 21.883 22.172 22.195 22.318 (17.044)
Table 2: PickScore training until reward collapses. Smallest reward gaps and best sampler performances locate at the large training steps $T=1200,1400$.
@@
$T=\bm{0}$ 100 200  $T=\bm{0}$ 100 200
Mean 0.382 (0.539) 0.668 (0.805) 0.915 (1.031)    0.347 (0.470) 0.763 (0.872) 1.086 (1.174)
Std 0.814 (0.799) 0.775 (0.721) 0.709 (0.652)    1.030 (0.997) 0.909 (0.868) 0.795 (0.727)
Gap 0.146 0.138 0.116  0.115 0.110 0.106
Table 3: Performance Comparison between prompts of different complexity. More complicated prompts yields faster fine-tuning improvements, larger in-group variances, and smaller SDE-ODE reward gaps.
@@ 5.2 MixGRPO
We use FLUX.1 (Black Forest Labs, 2024) as the base model, and fine-tune it with MixGRPO, which is a sliding-window sampler that alternates between ODE and SDE schemes for 25 training steps in total. Training is carried out with multiple rewards combined using equal weights, while evaluation is reported on ImageReward and HPSClip.
$\bullet$ Bounded Reward Gap Under High Stochasticity: With $\eta=1.2$, ODE sampling in MixGRPO consistently outperforms the mixed SDE–ODE scheme, in contrast to the results from DDPO. As shown in the top-left and bottom-left panels of Figure 4, the reward gap steadily diminishes during training and converges to zero.
$\bullet$ Quality Improvement: The middle and right panels of Figure 4 further shows that the ODE sampler performs consistently better on training prompts. For example, in the left pair of Figure 5, the generation with $\eta=1.2$ correctly aligns with the “trapped inside” prompt instruction, whereas the generation with $\eta=1.0$ fails to do so.
Figure 4: Bounded reward gap (left column) and performance improvement (middle, right column) for MixGRPO
@@ 6 Conclusion and Further Directions
Sampling stochasticity in diffusion RLHF plays an important role yet remains underexamined. Here we present an explicit solution for generalized VE/VP diffusion models, together with extensive experiments, demonstrating that “high stochasticity in training samples, no stochasticity in generation” is both theoretically sound and practically advantageous. Our results open a broader space for tuning stochasticity hyperparameters, enabling more robust and diversified diffusion post-training. Future work will quantify the “reward gap” in video and multimodal generation and develop clearer theoretical insights into stochasticity as a drifting force toward the target distribution.
@@ Appendix B Useful Propositions
A classical distance result on two Gaussian distributions (Hershey & Olsen, 2007) states:
@@ Appendix C Proofs for Section 3
Since $\bm{\mu}_{i}\perp\bm{r}$, $\mathbb{E}[Y_{\parallel}^{\theta}(0)]=0$. Also, the backward dynamic injects a scaled multiple of $\bm{I}_{d}$ noise, so the coordinate-wise dynamics are independent. Therefore, we are able to analyze the $\text{Span}(\bm{r})$ subspace via separating each dimension $\bm{e}_{k}\in\text{Span}(\bm{r})$, in which $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},SDE})\right\}$ and $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},ODE})\right\}$ follows a similar controlled motion as discussed in Theorem 3.1 and 3.2. Moreover, the score function can be bounded by a constant of $\max_{i}{\sigma_{i}},\min_{i}^{-1}{\sigma_{i}}$ (Liang et al., 2025).
@@ Appendix D Proof for Section 4
(Gronwall’s) If $u^{\prime}(t)\leq\alpha(t)u(t)+\beta(t)$ and in addition satisfies the $\kappa$-strict negativity $\alpha(t)\leq-\kappa<0$ for some positive constant $\kappa$,
$u(T)\leq e^{-\kappa T}u(0)+\int_{0}^{T}e^{-\kappa(T-s)}\beta(s)ds$ (17)
Proof.
The classical Gronwall’s theorem (Evans, 2010) gives
$u(T)\leq e^{\int_{0}^{T}\alpha(s)ds}\left(u(0)+\int_{0}^{T}e^{-\alpha(s)}\beta(s)ds\right).$
When $\alpha(t)\leq-\kappa$, $e^{\alpha(t)}\leq e^{-\kappa t}$ and the proposition follows naturally from integration. ∎
@@
$u(T)\leq\int_{0}^{T}e^{-\kappa(T-s)}\beta(s)ds.$ (18)
@@
$=2\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},-f(t,Y_{t}^{SDE})+f(t,Y_{t}^{ODE})\,\big\rangle$
$\qquad+g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{ODE})-s(t,Y_{t}^{SDE}))\,\big\rangle$
$\qquad-\eta^{2}g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{SDE})+\theta(t)\,\big\rangle+\eta^{2}g^{2}(t)$
$\leq 2\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},-f(t,Y_{t}^{SDE})+f(t,Y_{t}^{ODE})\,\big\rangle$
$\qquad+||g||_{\infty}^{2}\cdot\mathbb{E}\big\langle\,Y_{t}^{ODE}-Y_{t}^{SDE},s(t,Y_{t}^{ODE})-s(t,Y_{t}^{SDE}))\,\big\rangle$
[…]
By Condition 2 and 3,
[…]
Together with Young’s Inequality,

### v2

@@ 3.1 VE with a Gaussian target
$dY_{t}^{\text{REF}}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}Y_{t}^{\text{REF}}dt+\eta\sqrt{2(T-t)}dB_{t}.$
Next we set the reward function $r(x)=-(x-1)^{2}$, so the goal of fine-tuning is to drive the sample towards mean $1$. We also specify the fine-tuned SDE and ODE as defined in equation 9 by
$dY_{t}^{\text{SDE}}=-\frac{(1+\eta^{2})(T-t)}{1+(T-t)^{2}}(Y_{t}^{\text{SDE}}+\theta_{\eta}^{*})dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{\text{ODE}}=-\frac{T-t}{1+(T-t)^{2}}(Y_{t}^{\text{ODE}}+\theta_{\eta}^{*})dt,$
@@
For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-(1+T^{2})^{-\frac{1+\eta^{2}}{2}}\right)\right]^{-1}$. Moreover, for $T\geq 1$
@@ 3.2 VP with a Gaussian target
Under the same setup as equation 9 and the VE case, we have:
$dY_{t}^{\text{REF}}=-\eta^{2}(T-t)Y_{t}^{\text{REF}}dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{\text{SDE}}=-\eta^{2}(T-t)Y_{t}^{\text{SDE}}\,dt-(1+\eta^{2})(T-t)\theta^{*}_{\eta}(t)\,dt+\eta\sqrt{2(T-t)}dB_{t},$
$dY_{t}^{\text{ODE}}=-(T-t)\theta_{\eta}^{*}(t)dt,$
@@
Consider the VP model (equation 13), with the reward function $r(x)=-(x-1)^{2}$. For $\eta>0$, we have $\theta^{*}_{\eta}=-\left[(1+\frac{\beta}{2})\left(1-e^{-\frac{(1+\eta^{2})T^{2}}{2}}\right)\right]^{-1}$. Moreover, for $T\geq 1$
@@ 3.3 VE/VP with a mixture Gaussian target
For all these examples, the reward gap $\Delta_{\eta}\downarrow 0$ (independent of $\eta$), as $T\to\infty$. Such a phenomenon will also be observed in fine-tuning T2I models with more complex rewards.
@@ 4 Non-vacuous bound on the reward gap
[…]
@@
The following conditions hold for all $y,y_{1},y_{2}$ along inference trajectories,
@@
Condition 1 (Lipschitz of scores) and 2 ($L^{2}$ bound) are standard in stability analysis. Also, there always exists a continuous-time VP/VE SDE on $[0,1]$ whose discretization (ddim) matches the large-scale T2I models’ (i.e. Stable Diffusion and SDXL) training schedule. In addition, models (FLUX) based on rectified flow inherently uses a normalized time interval. Therefore, for large $\eta$, the following theorem yields $O(\eta^{2})$ bound on $W_{2}$ distance, see proof in Appendix D.1.
@@
If Assumption 4.1 hold with $T=1$, $W_{2}(Y_{T}^{\text{ODE}},Y_{T}^{\text{SDE}})\leq C\cdot\max\{\eta,\eta^{2}\}\cdot\sqrt{1+A}$, where $C$ only depends on $g||_{\infty}$ and $L$.
Proposition 4.1.
Let Assumption 4.1 hold with $L_{r}$-Lipschitz rewards $r$, $\Delta_{\eta}\leq L_{r}\cdot C\cdot\max\{\eta,\eta^{2}\}\cdot\sqrt{1+A}$, where $C$ only depends on $g||_{\infty}$ and $L$.
An improved $O(\eta)$ bound relies on contractive coefficients (see Appendix D.2). This contractivity can arise from a contractive drift $f$ (TZ24). Alternatively, even though the backward dynamics of classical VP models are expansive, a strongly log-concave terminal distribution (GNZ23) still satisfies the contractive condition (see Assumption D.1), which can hold if the conditional terminal distribution for a fixed prompt $\bm{c}$ is approximately unimodal.
Quadratic and linear growth rate on $\eta$ is consistent with our empirical observation: T2I fine-tuning preserves quality for ODE inference; rewards only deteriorate under very large $\eta$ (see Table 3).
Remark.
Discretization errors can be incorporated into the $W_{2}$ bound as an additional term $\epsilon_{d}(h)$ depending on the time–step size $h$ (gaussianmixture), with $\epsilon_{d}(h)\to 0$ as $h\to 0$. Under Assumption 4.1, Euler–Maruyama yields such a vanishing term; see Appendix D.3 for details.
@@ 5 Numerical experiments
In this section, we fine-tune large-scale T2I models under large $\eta$ and examine its reward gap $\Delta_{\eta}$ with RLHF algorithms DDPO and MixGRPO (see Section 2.3). Preference rewards for fine-tuning DDPO include the LAION aesthetic (aesthetic), HPS_v2.1 (hpsv2), PickScore (pickscore), and ImageReward (imagereward). The rewards for fine-tuning MixGRPO include HPS CLIP, PickScore, ImageReward, and Unified Reward (unifiedreward).
According to equation 4 and equation 5, we use a high stochasticity $\eta\geq 1.0$ under gDDIM scheme (see Section 2.2 and Appendix E) to generate robust training samples $\{Y_{T}^{\text{SDE}}\}$ and compare them with deterministic inference samples $\{Y_{T}^{\text{ODE}}\}$ following equation 9. The stochasticity scale $\eta$ controls the noise level of backward dynamics.
Remark.
In each experiment, the training stochasticity $\eta$ and the number of training steps $N$ are always specified. Choices of other hyperparameters are detailed in Appendix F.
@@ 5.1 DDPO
We use Stable Diffusion v1.5 (stabledif) as the base model, and fine-tune it with DDPO. During training, we adopt ImageReward and PickScore as preference rewards, while Aesthetic and HPS_v2 are included as additional evaluation metrics. Figure 7 (and Appendix H) provides representative generations from the fine-tuned models. Our main observations are summarized as follows:
$\bullet$ High Stochasticity Benefits Moderate Training Steps. We compare fine-tuning under ImageReward and PickScore at $\eta\in\{1.0,1.2,1.5\}$ for $N=200$ steps. As shown in Table 1, $\eta=1.2$ under ImageReward achieves the best in-domain and out-of-domain performance, while PickScore’s performances depend on evaluation metrics.
$\bullet$ Decreasing Reward Gap with Quality Improvement. To study the reward gaps for smaller or larger $N$, we experiment under PickScore with $\eta=1.2$ and calculate SDE–ODE reward differences (here we report the SDE over ODE performance) under multiple preference functions every 200 steps until reward collapse. As shown in Table 2, the gap decreases as image quality improves for both samplers.
$\bullet$ Richer Prompt Contents Reduce Reward Gap. We compare performances with animal versus more comprehensive prompts (see Appendix H.4) under ImageReward with $\eta=1.2$. As shown in Table 3, complex prompts generate higher post-tuning rewards and higher in-group variance in post-training. Moreover, their richer instructions reduce the SDE–ODE reward gap.
Figure 3: Evolution of reward gap (the reward difference between SDE sampling and ODE sampling) during DDPO training under PickScore fine-tuning with stochasticity $\eta\in\{1.0,1.2,1.5\}$. The gap for multiple rewards remains small as training step NN progresses.
$N=\bm{0}$ 200 400 600 800 1000 1200 1400 (1476)
@@
SDE Reward 20.82 21.31 21.65 21.90 22.07 22.27 22.31 22.42 (18.92)
ODE Reward 20.73 21.17 21.50 21.78 21.88 22.17 22.20 22.32 (17.04)
Table 2: PickScore training with $\eta=1.2$ until reward collapses. Smallest reward gaps and best sampler performances locate at the large training steps $N=1200,1400$.
@@
$N=\bm{0}$ 100 200  $N=\bm{0}$ 100 200
Mean 0.38 (0.540) 0.67 (0.81) 0.92 (1.03)    0.35 (0.47) 0.76 (0.87) 1.09 (1.17)
Std 0.81 (0.80) 0.78 (0.72) 0.71 (0.65)    1.03 (1.00) 0.91 (0.87) 0.80 (0.73)
$\Delta_{\eta=1.2}$ 0.15 0.14 0.12  0.12 0.11 0.11
Table 3: Performance Comparison between prompts of different complexity with $\eta=1.2$. More complicated prompts yields faster fine-tuning improvements, larger in-group variances, and smaller SDE-ODE reward gaps.
@@ 5.2 MixGRPO
We use FLUX.1 (flux) as the base model, and fine-tune it with MixGRPO, which is a sliding-window sampler that alternates between ODE and SDE schemes for 25 training steps in total. Training is carried out with multiple rewards combined using equal weights, while evaluation is reported on ImageReward and HPSClip.
$\bullet$ Bounded Reward Gap Under High Stochasticity. As shown in Figure 4, $\Delta_{\eta}$ converges to zero for both HPS Clip and ImageReward. Also, with $\eta=1.2$ and ImageReward metric, ODE sampling in MixGRPO consistently outperforms the mixed SDE–ODE scheme, in contrast to the results from DDPO.
Figure 4: Bounded reward gap for MixGRPO, displayed with 7-step moving average. Here a positive score difference means the SDE component achieves higher reward than the ODE component.
$\bullet$ T2I Quality Improvement. The upper (evaluated with ImageReward) and lower panels (evaluated with HPS Clip) of Figure 5 further show that samplers trained with larger stochasticity ($\eta=1.2$) perform consistently better than standard DDIM stochasticity ($\eta=1.0$) on training prompts. For example, in Figure 6, the generation with $\eta=1.2$ correctly aligns with the “trapped inside” prompt instruction, whereas the generation with $\eta=1.0$ fails to do so.
Figure 5: Performance improvement for MixGRPO, displayed with 7-step moving average.
@@ 6 Conclusion and Further Directions
This work clarifies the tension between stochastic SDE training and deterministic ODE inference in diffusion RLHF. By proving a bounded reward gap $\Delta_{\eta}$ and empirically showing that higher training stochasticity (e.g., $\eta=1.2$) improves deterministic image quality, we provide theoretical support for “training with SDE, inference with ODE”. Future work includes quantifying how distribution shift and the choice of reward function separately affect reward gaps.
Acknowledgments
Wenpin Tang and Jiayuan Sheng are supported by NSF grant DMS-2206038. Wenpin Tang acknowledges support by the Columbia Innovation Hub grant, and the Tang Family Assistant Professorship. The works of Haoxian Chen, Jiayuan Sheng, Hanyang Zhao, and David Yao are part of a ColumbiaCityU/HK collaborative project that is supported by InnoHK Initiative, The Government of the HKSAR and the AIFT Lab.
@@ Appendix B Useful Propositions
A classical distance result on two Gaussian distributions (4218101) states:
@@ Appendix C Proofs for Section 3
Since $\bm{\mu}_{i}\perp\bm{r}$, $\mathbb{E}[Y_{\parallel}^{\theta}(0)]=0$. Also, the backward dynamic injects a scaled multiple of $\bm{I}_{d}$ noise, so the coordinate-wise dynamics are independent. Therefore, we are able to analyze the $\text{Span}(\bm{r})$ subspace via separating each dimension $\bm{e}_{k}\in\text{Span}(\bm{r})$, in which $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},SDE})\right\}$ and $\left\{\text{Proj}_{\bm{e}_{k}}(Y_{\parallel}^{\theta_{\parallel}^{*},ODE})\right\}$ follows a similar controlled motion as discussed in Theorem 3.1 and 3.2. Moreover, the score function can be bounded by a constant of $\max_{i}{\sigma_{i}},\min_{i}^{-1}{\sigma_{i}}$ (gaussianmixture).
@@ Appendix D Proof for Section 4
(Gronwall’s) Suppose integrable functions $u,\alpha,\beta:\mathbb{[}0,T]\to\mathbb{R}$ satisfies $u^{\prime}(t)\leq\alpha(t)u(t)+\beta(t)$ ,
$u(T)\leq e^{\int_{0}^{T}\alpha(s)ds}\left(u(0)+\int_{0}^{T}e^{-\int_{0}^{s}\alpha(\tau)d\tau}\beta(s)ds\right).$ (17)
A proof can be found in (evans).
D.1 Proof of Theorem 4.1
@@
$u(T)\leq e^{\alpha T}\left(\int_{0}^{T}\beta\cdot e^{-\alpha s}\,ds\right)=\frac{(e^{\alpha T}-1)\cdot\beta}{\alpha}.$ (18)
@@
$=2\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},-f(t,Y_{t}^{\text{ODE}})+f(t,Y_{t}^{SDE})\,\big\rangle$
$\qquad\;+g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}})\,\big\rangle$
$\qquad\;-g^{2}(t)\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},\eta^{2}\cdot s_{\theta}(t,Y_{t}^{\text{SDE}})\,\big\rangle+\eta^{2}g^{2}(t).$
According to (TZ24tut),
$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},-f(t,Y_{t}^{\text{ODE}})+f(t,Y_{t}^{\text{SDE}})\,\big\rangle=\left\{\begin{array}[]{ll}0&\text{for VE},\\ g^{2}(t)u(t)/2&\text{for VP}.\end{array}\right$.
By Condition 1,
$\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},s_{\theta}(t,Y_{t}^{\text{ODE}})-s_{\theta}(t,Y_{t}^{\text{SDE}}))\,\big\rangle$ $\leq\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},L\cdot\left(Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}}\right))\,\big\rangle$
$=L\cdot u(t).$
By Condition 2,
$\left|\mathbb{E}\big\langle\,Y_{t}^{\text{ODE}}-Y_{t}^{\text{SDE}},\eta^{2}\cdot s_{\theta}(t,Y_{t}^{\text{SDE}})\,\big\rangle\right|$ $\leq u(t)+\frac{\eta^{4}}{4}\cdot\mathbb{E}[||s_{\theta}(t,Y_{t}^{\text{SDE}})||^{2}$
$=u(t)+\frac{\eta^{4}}{4}\cdot A.$
Together, by absorbing $g||_{\infty}$ into an $\eta$-free constant $C_{\text{scheme}}$,
@@
And we apply equation 18 with $T=1$:
$u(T)$ $\leq\frac{(e^{\alpha T}-1)\cdot\beta}{\alpha}$
$\leq\frac{\exp({C_{\text{scheme}}\cdot L})}{C_{\text{scheme}}\cdot L}\cdot\max\{\eta^{4},\eta^{2}\}\cdot\left(\frac{A}{4}+1\right).$
Finally, the $W_{2}$ distance can be bounded in terms of this $L^{2}$–distance via the chosen coupling:
$W_{2}\big(\mathcal{L}(Y_{T}^{\mathrm{ODE}}),\mathcal{L}(Y_{T}^{\mathrm{SDE}})\big)\;\leq\;\Big(\mathbb{E}\big[\|Y_{T}^{\mathrm{ODE}}-Y_{T}^{\mathrm{SDE}}\|^{2}\big]\Big)^{1/2}\;\leq\;C(\|g\|_{\infty},L)\,\max\{\eta,\eta^{2}\}\,\sqrt{1+A}.$
D.2 Proof for strongly log-concave distributions
Assumption D.1.
(Strong log-concavity): Exists $\kappa>1$ such that for all $t,y_{1},y_{2}$: